find the present value of $7,00 due in 4 years if the money is worth:
a. 6% compunded smei-anually
b. 6% compunded anually
c. 6% compunded quarterly

Answer by rfer(16322)

P=A/(1+r/n)^nt
P=7000/(1+0.06/2)^2*4
P=7000/(1.03)^8
P=7000/1.26677
P=$5,525.86
--------------------
You do the others.